OpenStudy (anonymous):

What is the molar concentration of sodium ions in a solution when 184.82 grams of sodium carbonate is added to water to produce 2.82 liters of solution?(assuming NaCO3 is a strong electrolyte)

OpenStudy (anonymous):

Na2CO3 -> 2Na^+ + CO3^2- n(Na2CO3) = m(Na2CO3)/M(Na2CO3) n(Na+) / n(Na2CO3) = 2/1 = 2 n(Na+) = 2*n(Na2CO3) c=n/V c(Na+) = n(Na+) / V(solution)